OpenStudy (anonymous):

The magnitude and direction of vectors u and v are given. Find vector w's polar coordinates. u: magnitude 140, bearing 160° v: magnitude 200, bearing 290° w: u + v A. w: (282, 141°) B. w: (23603, 226°) C. w: (153.6, 246°) D. w: (22, 182°)

OpenStudy (anonymous):

@satellite73 Can you help?

OpenStudy (anonymous):

u: magnitude 140, bearing 160° v: magnitude 200, bearing 290° find u + v First convert polar (r,θ) to rectangular (x,y) x = r cos(θ) y = r sin(θ) u: x = 140 cos 160° and y = 140 sin 160° x = -131.557 and y = 47.883 v: x = 200 cos 290° and y = 200 sin 290° x = 68.404 and y = -187.939 add x and y components x = -131.557 + 68.404 = -63.153 y = 47.883 + -187.939 = -140.056 convert rectangular coordinates back to polar coordinates r = √-63.153² + -140.056² r = 153.6 θ = 270 - arctan(63.153/140.056) θ = 245.73 w: (153.6, 245.73°)

OpenStudy (anonymous):

I can try, but i have to warn you im not that good at these